DC circuits Begin Date: 2/14/2017 12:00:00 AM Due Date: 2/20/2017 5:00:00 PM End Date: 3/9/2017 12:00:00 AM (13%) Problem 2: An automobile starter motor has an equivalent resistance of 0.075 2 and is supplied by a 12.0 V battery which has a 0.045 S2 internal resistance A 33% Part (a) What is the current to the motor in A? Grade Summary Deductions Potential 100% sino cos0 tan0 It 7 8 9 Submissions HOME Attempts remaining: 7 4 5 6 asino acos0 cotan 4% per attempt) detailed view atan0 acotan0 sinh0 1 2 3 cosh0 tanh0 cotanh0 Degrees Radians BACKSPACE Submit Hint Feedback I give up! Hints: 49 deduction per hint. Hints remaining: 2 Feedback: 5% deduction per feedback. A 33% Part (b) What voltage is applied to it in V? A 33% Part (c) What power is supplied to the motor in kW? All content 2017 Expert TA, LLCI need help with all parts
Answer #1

Equivalent resistance, Rm = 0.075 Ohm.

Internal resistance, r = 0.045 ohm.

Supply voltage, V = 12 V

Part (a):

Internal resistance of battery is in series with the motor winding resistance.

So, equivalent resistance, R = Rm + r = 0.075 + 0.045 = 0.120 V

So, current in the motor = V/R = 12/0.12 = 100 A

Part (b):

Applied voltage, E = I*Rm = 100*0.075 = 7.5 V

Part (c):

Power supplied to the motor, W = I^2*Rm = 100^2*0.075 = 750 W